# Trustworthy Answers, Messier Data: Bridging the Gap in Low-Resource Retrieval-Augmented Generation for Domain Expert Systems

### v2

@@
Abstract
RAG has become a key technique for enhancing LLMs by reducing hallucinations, especially in domain expert systems where LLMs may lack sufficient inherent knowledge. However, developing these systems in low-resource settings introduces several challenges: (1) handling heterogeneous data sources, (2) optimizing retrieval phase for trustworthy answers, and (3) evaluating generated answers across diverse aspects. To address these, we introduce a data generation pipeline that transforms raw multi-modal data into structured corpus and Q&A pairs, an advanced re-ranking phase improving retrieval precision, and a reference matching algorithm enhancing answer traceability. Applied to the automotive engineering domain, our system improves factual correctness (+1.94), informativeness (+1.16), and helpfulness (+1.67) over a non-RAG baseline, based on a 1-5 scale by an LLM judge. These results highlight the effectiveness of our approach across distinct aspects, with strong answer grounding and transparency.
Trustworthy Answers, Messier Data: Bridging the Gap in Low-Resource Retrieval-Augmented Generation for Domain Expert Systems
Nayoung Choi1, Grace Byun1, Andrew Chung1, Ellie S. Paek1, Shinsun Lee††thanks: Work done while at Emory University as a visiting scholar.1,2, Jinho D. Choi1 1Emory University, 2Hyundai Motor Company {nayoung.choi, grace.byun, andrew.chung, ellie.paek, shinsun.lee, jinho.choi}@emory.edu
1 Introduction
Retrieval-Augmented Generation (RAG) has shown potential in reducing hallucinations and providing up-to-date knowledge in Large Language Models (LLMs). This success has grown interest in domain expert RAG-Question Answering (QA) systems to meet specialized knowledge needs. While previous studies (Han et al. 2024; Siriwardhana et al. 2023; Mao et al. 2024) have proposed general methods for adapting RAG models to domain knowledge bases—such as syntactic QA pair generation or model fine-tuning—they face several challenges in low-resource domains.
@@ 1 Introduction
In practical settings, available data sources in low-resource domains are often presented in heterogeneous formats and exhibit an unstructured nature, making their direct integration into RAG system development challenging Hong et al. (2024). General models may not have enough inherent knowledge in low-resource domains Zhao et al. (2024), making fine-tuning essential to adapt the model to specific knowledge requirements. However, the lack of structured data for training further complicates this process. In addition, data privacy and security concerns restrict the full utilization of API-based LLMs (Achiam et al. 2023; Anthropic 2024) within RAG systems, necessitating the use of open-source LLMs (Yang et al. 2024; Touvron et al. 2023).
The retrieval phase is another key aspect of domain expert RAG systems, as referencing accurate documents is essential for generating reliable answers. However, research on improving ranking in the retrieval phase or tracing the documents referenced to generate the answer remains limited. Most domain expert RAG frameworks rely on a single-stage retrieval process, with few exploring multi-stage approaches (Nogueira et al. 2019; Nogueira et al. 2020; Karpukhin et al. 2020), such as retrieval followed by re-ranking—a method widely used in Information Retrieval (IR)—which can help ensure the use of the most relevant references.
The evaluation of RAG systems is also an area that has not been fully addressed. Many studies continue to rely on gold answer similarity metrics, such as overlapping words between the generated answer and the ground truth, which inadequately capture critical dimensions like faithfulness, coherence, and contextual relevance. Recently, the LLM-as-a-judge framework Zheng et al. (2023) has gained attention as a qualitative alternative. However, these methods often overlook diverse evaluation aspects, and a standardized framework for assessing RAG systems has yet to emerge.
In this work, we address three key challenges and present the RAG development pipeline, demonstrating its application in the automotive engineering domain, with a specific focus on QA for vehicle crash collision tests. First, we present a data generation pipeline (Section 3.1), leveraging diverse formats of internal documents from an automobile company. Second, we incorporate an advanced re-ranking phase (Section 3.2) and a reference-matching algorithm (Section 3.4), not only enhancing the accuracy of the final answers but also improving traceability by tagging sources for each segment of the answer. Third, we evaluate the answers obtained from our generation model (Section 3.3) from multiple perspectives, emphasizing their qualitative aspects. Our ultimate goal is a fully local system, ensuring data privacy and independence from external servers. The overall workflow of our RAG system is shown in Figure 1, and our key contributions are summarized as follows:
We propose a data generation pipeline that transforms multi-modal raw data into a structured corpus and high-quality Q&A pairs.
We integrate re-ranking and reference matching to enhance retrieval precision and answer traceability, ensuring a reliable RAG system.111The full RAG pipeline code will be publicly available: https://github.com/emorynlp/MultimodalRAG
We assess the final answer from diverse qualitative perspectives, evaluating each along distinct, non-overlapping dimensions.
2 Related Work
@@
Data Processing RAG-Studio (Mao et al. 2024) employs synthetic data generation for in-domain adaptation, reducing reliance on costly human-labeled datasets. However, it assumes access to well-structured data, limiting its applicability in scenarios with unstructured raw data. To bridge this gap, Hong et al. (2024) tackle challenges with real-world formats (e.g., DOC, HWP), proposing a chunking method that converts documents to HTML, retaining structural information in low-resource settings. Meanwhile, Guan et al. (2024) address the issue of short and noisy e-commerce datasets in RAG system development by building a new dataset from a raw corpus crawled from an e-commerce website, providing a richer resource.
Retrieval in RAG Wang et al. (2024) highlight the importance of re-ranking modules in RAG systems to enhance the relevance of the retrieved documents. Similarly, Zhao et al. (2024) demonstrate that the ranking position of the gold document in the retrieval phase plays a significant role in determining the quality of the final answer. Given these insights, optimizing the retrieval phase is crucial for obtaining accurate context, particularly in specialized, low-resource domains where LLMs lack sufficient inherent knowledge Beauchemin et al. (2024). Despite these findings, the analysis of the effectiveness and applicability of re-rankers in domain expert RAG systems remains underexplored.
@@
RAG Evaluation The evaluation of RAG systems Yu et al. (2024) has relied on text similarity-based metrics such as BLEU Papineni et al. (2002), ROUGE Lin (2004), and EM Rajpurkar et al. (2016). These metrics provide a baseline but overlook valid answer diversity and qualitative aspects like factual consistency and relevance. Recent advancements are made to utilize LLM-as-a-judge Zheng et al. (2023) to evaluate qualitative dimensions. Han et al. (2024) propose a pairwise preference comparison framework considering helpfulness and truthfulness, while Saad-Falcon et al. (2024) assess context relevance, answer faithfulness, and query relevance, addressing hallucination issues. However, a standardized framework for RAG evaluation remains a challenge.
3 Approach
3.1 Data Generation & Processing
We build our dataset using two distinct sources: (1) internal reports on vehicle crash collision tests in presentation slide format from an automotive company in Korea, and (2) the textbook Crash Safety of Passenger Vehicles Mizuno (2016). Vehicle crash collision tests are exceptionally valuable, as each test incurs substantial costs and produces detailed reports that are typically confidential and difficult to access publicly. Figure 2 illustrates the process of extracting, analyzing, and converting slides and textbook PDFs into structured Markdown text and Q&A sets, leveraging Python scripts and the Claude LLM222We restrict the use of API-based LLMs for initial data processing; this limitation is discussed in Section 8. Anthropic (2024). Table 1 summarizes the data statistics by source.
The original slides often contain tables, graphs, and images, which are simplified into plain text descriptions during data processing. To evaluate performance on questions referencing these elements, we sample 143 multi-modal chunks and generated 1,861 targeted questions. Evaluation results are provided in Section 4.3.1.
@@ 3.1 Data Generation & Processing
Total | 5,625 | 68,603 |
Additionally, we extract and prepend headers summarizing each report, including the Test Name, Region, State, and Purpose, which significantly improve the retrieval phase (See Section 4.3). Domain experts are responsible for the entire process, including prompt engineering and reviewing intermediate and final outputs. All prompts and an example of a generated Q&A pair are shown in A.2.1.
3.2 Retrieval & Ranking
We employ a Dual-Encoder as our retriever, fine-tuning it on our training data. For a question , the model retrieves the top chunks from chunks , based on the [CLS] token embedding similarity between and each chunk , as follows:
Our re-ranker is a point-wise Cross-Encoder, trained on a classification task Nogueira et al. (2019) that takes a single pair as input and returns a scalar relevance score. It re-ranks , obtained from the retrieval stage, to extract , which will be passed to the generation model. The process is formalized as follows:
To optimize , we apply Token Selection (TS) + Term Control Layer (TCL) training method from RRADistill Choi et al. (2024). It effectively integrates the importance of the general semantic and specific query terms during the training process.
TS+TCL are used during training only, while inference uses the standard Cross-Encoder approach only with . The target label is binary: for relevant pairs and for irrelevant pairs. Negative sampling retrieves the top 10 chunks from the train data for each query , excluding the gold chunk, and randomly selects three from the rest.
3.3 Answer Generation
@@
We fine-tune open-source LLMs on our training dataset, which consists of pairs derived from our data generation pipeline (Section 3.1). During fine-tuning, the question and the answer are concatenated into a single sequence , and the model is trained in an auto-regressive manner to predict the next token.
3.4 Reference Matching Algorithm
We propose a reference matching algorithm that segments generated answers and links them to relevant references using a Dynamic Programming (DP) approach Bellman (1954). Algorithm 1 outlines the detailed procedure.
@@ 3.4 Reference Matching Algorithm
In Step 1, the algorithm computes the scores for all possible sentence subsequences against the top-k chunks , using our re-ranker (Section 3.2). In Step 2, it selects the optimal segment-chunk combinations, updating scores and recording the best choices for each ending sentence (e.g., if the answer consists of five sentences like , the choices might be ). Finally, in Step 3, backtracking is performed to retrieve and output the best matches from the choices (e.g., ).
4 Experiment 1: Retrieval & Ranking
4.1 Models
For retriever training, we use BGE-M3 Chen et al. (2024) as the backbone, a multilingual Encoder capable of processing Korean, and fine-tune it on our training dataset with the publicly available code333https://github.com/FlagOpen/FlagEmbedding. For re-ranker training, we initialize the weights using the fine-tuned retriever. Further experimental details are in A.1.2.
4.2 Evaluation
@@
We evaluate the retriever and re-ranker using Mean Average Precision (MAP@k) and Success@k Manning et al. (2008). MAP@k measures the ranking quality by averaging precision over relevant results up to rank , while Success@k indicates the proportion of queries with at least one relevant result in the top . The high Success@k score indicates that relevant chunks are retrieved, while the improved MAP@k highlights better ranking of those retrieved chunks. The evaluation is conducted using test set questions, but the chunk pool for retrieval included all splits (training, validation, and test) to ensure sufficient data and avoid biases from the limited test set size.
4.3 Result
@@
Table 2 shows that prepending a header to each chunk—adding only a small number of tokens (see Table 9)—significantly enhances retrieval performance. Fine-tuning the BGE-M3 model on Ver.1 (BGE-FT) further improves results notably, demonstrating the importance of task-specific model adaptation to optimize performance. Table 3 shows that the re-ranking phase notably enhances the precision of retrieved information. The results indicate that ranking performance is optimal when the number of retrieved chunks is limited to 10. Despite the constraints of retriever failure, this improves retrieval results by about 4% in terms of MAP@1. In addition, incorporating TS+TCL during training boosts ranking performance (See A.3.1).
Model | Train | Test | MAP@1 | MAP@5 | MAP@10 |
BGE- Vanilla | N/A | Ver.0 | 0.1978 | 0.2672 | 0.2766 |
Ver.1 | 0.2991 | 0.3978 | 0.4091 | ||
@@
MAP@1 | MAP@5 | MAP@10 | Success@5 | |||||||||
10 |
20 |
30 |
4.3.1 Multi-modal Specific Questions
Table 4 summarizes the performance on questions requiring information from tables, graphs, and images. Compared to general questions (Section 4.3), the results are notably lower, with the retriever facing significant challenges. Even at = 30, Success@k only reaches 83.61%, underscoring a major bottleneck in the retrieval phase. Despite these challenges, the re-ranker proved effective, delivering a substantial performance improvement of approximately 15% in MAP@1 at = 20. These results underscore the need for further advancements to better handle multi-modal questions.

### v5

# Reference-Aligned Retrieval-Augmented Question Answering over Heterogeneous Proprietary Documents

@@
Abstract.
Proprietary corporate documents contain rich domain-specific knowledge, but their overwhelming volume and disorganized structure make it difficult even for employees to access the right information when needed. For example, in the automotive industry, vehicle crash-collision tests—each costing hundreds of thousands of dollars—produce highly detailed documentation. However, retrieving relevant content during decision-making remains time-consuming due to the scale and complexity of the material. While Retrieval-Augmented Generation (RAG)-based Question Answering (QA) systems offer a promising solution, building an internal RAG-QA system poses several challenges: (1) handling heterogeneous multi-modal data sources, (2) preserving data confidentiality, and (3) enabling traceability between each piece of information in the generated answer and its original source document. To address these, we propose a RAG-QA framework for internal enterprise use, consisting of: (1) a data pipeline that converts raw multi-modal documents into a structured corpus and QA pairs, (2) a fully on-premise, privacy-preserving architecture, and (3) a lightweight reference matcher that links answer segments to supporting content. Applied to the automotive domain, our system improves factual correctness (+1.79, +1.94), informativeness (+1.33, +1.16), and helpfulness (+1.08, +1.67) over a non-RAG baseline, based on 1–5 scale ratings from both human and LLM judge. The system was deployed internally for pilot testing and received positive feedback from employees.
1. Introduction
@@
Retrieval-Augmented Generation (RAG) has shown potential in reducing hallucinations and providing up-to-date knowledge in Large Language Models (LLMs). This success has grown interest in domain expert RAG-Question Answering (QA) systems to meet specialized knowledge needs. While previous studies (Han et al., 2024) (Siriwardhana et al., 2023) (Mao et al., 2024) have proposed general methods for adapting RAG models to domain knowledge bases—such as syntactic QA pair generation or model fine-tuning—they face several challenges for internal use in enterprise settings. Internal corporate documents are often presented in heterogeneous formats and exhibit an unstructured nature, making their direct integration into RAG system development challenging (Hong et al., 2024). Moreover, general-purpose LLMs often lack sufficient knowledge of specialized corporate domains (Zhao et al., 2024), making fine-tuning essential. However, this process is further complicated by the scarcity of structured training data. In addition, data privacy and security concerns restrict the full utilization of API-based LLMs (Achiam et al., 2023) (Anthropic, 2024) within RAG systems, necessitating the use of open-source alternatives (Touvron et al., 2023) (Yang et al., 2024). These limitations highlight the need for a RAG-QA pipeline tailored for internal enterprise use, one that can process heterogeneous documents while ensuring data privacy.
The automotive industry is one of the domains where an internal RAG-QA pipeline is particularly needed, due to both the scale and sensitivity of engineering data. For example, a single vehicle crash‑collision test 111A vehicle crash-collision test involves driving a car into a fixed barrier to measure occupant and pedestrian safety; each test costs hundreds of thousands of dollars. produces hundreds of pages of confidential artifacts, including design rationales, sensor logs, finite‑element analyses. Engineers review this documentation repeatedly when diagnosing failures, refining component designs, or compiling regulatory submissions, yet the material is fragmented across file shares and internal platforms, making retrieval slow. An internal RAG-QA system enables users to retrieve necessary information without manually navigating large volumes of documents, and to query LLMs without concerns about data confidentiality—ultimately reducing the time required for decision-making.
In this work, we address three key challenges of building a RAG-QA system for internal use and present an on-premise framework, demonstrating its application in a leading automotive company based in Korea. First, to handle heterogeneous and multi-modal documents, we present a data processing pipeline (Section 3.1) that converts raw inputs, such as meeting notes, test reports, and educational materials, into a structured corpus and the corresponding QA pairs. Second, to preserve data confidentiality, we propose an on-premise RAG system (Section 3.2) composed of domain-adapted retrieval, re-ranking, and generation components, none of which rely on external servers. Third, to support verifiable decision-making, we introduce a lightweight reference matching algorithm (Section 3.3) that links each answer segment to its supporting document, allowing users to trace the original sources and validate the information. The workflow of our RAG system is shown in Figure 1, and our key contributions are summarized as follows:
We present a data processing pipeline that converts heterogeneous multi-modal documents into Markdown text chunks and generates Q&A pairs to train the RAG system.
We develop a framework that domain-adapts the retrieval, re-ranking, and generation components of the RAG system and deploys them in an on-premise system.
We introduce a reference matching algorithm that uses dynamic programming to align answer segments with retrieved documents, enabling traceability to original sources.
2. Related Work
@@
Domain-specific RAG RAG-Studio (Mao et al., 2024) employs synthetic data generation for in-domain adaptation, reducing reliance on costly human-labeled datasets. However, it assumes access to well-structured data, limiting its applicability to real-world cases with unstructured raw inputs. To address such challenges, Hong et al. (2024) handle diverse formats (e.g., DOC, HWP) by converting documents to HTML while preserving structural cues, and Guan et al. (2024) mitigate the issue of short, noisy e-commerce data by constructing a richer dataset from raw crawled corpora. Building on these efforts, recent multimodal and enterprise-focused RAG systems—such as VisDoMRAG, which integrates visual and textual retrieval for visually rich documents (Suri et al., 2025), and eSapiens, which combines Text-to-SQL with RAG in enterprise settings (Shi et al., 2025)—further underscore the need for robust RAG methods capable of handling heterogeneous data.
@@
RAG Evaluation The evaluation of RAG systems (Yu et al., 2024) has relied on text similarity-based metrics such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and EM (Rajpurkar et al., 2016). These metrics provide a baseline but overlook valid answer diversity and qualitative aspects like factual consistency and relevance. Recent advancements are made to utilize LLM-as-a-judge (Zheng et al., 2023) to evaluate qualitative dimensions. Han et al. (2024) propose a pairwise preference comparison framework considering helpfulness and truthfulness, while Saad-Falcon et al. (2024) assess context relevance, answer faithfulness, and query relevance, addressing hallucination issues. However, a standardized framework for RAG evaluation remains a challenge, and the design of evaluation metrics should be tailored to the specific application context.
3. Approach
3.1. Data Generation & Processing
We use three different internal data sources from a Korean automotive company: (1) meeting notes, (2) test reports on vehicle crash-collision tests—both in presentation slide format—and (3) the educational text material Crash Safety of Passenger Vehicles (Mizuno, 2016). The first two sources contain a mix of multi-modal elements, including images, tables, and graphs. Figure 2 illustrates the process of extracting, analyzing and converting slides and textbook PDFs into structured Markdown text and Q&A sets, leveraging Python scripts and the Claude (Anthropic, 2024) Bedrock API 222Amazon Bedrock service ensures all prompts and responses remained within a secure enterprise-grade environment, aligning with our data privacy requirements.. The converted Markdown text is segmented into chunks—each slide is treated as one chunk for the slide sources, and each chapter as one chunk for the textbook. Table 1 presents the data statistics by source, while Table 2 shows the dataset split into training, validation, and test sets. Token length statistics for the training data are provided in Table 3.
@@ 3.1. Data Generation & Processing
| Total | 5,625 | 68,603 | ||
| Source | Data | Train | Val | Test | Total |
| Test Report | Chunk | 3,729 | 466 | 467 | 4,662 |
| Q&A Pair | 47,660 | 5,823 | 5,919 | 59,402 | |
| Meeting Note | Chunk | 705 | 88 | 89 | 882 |
| Q&A Pair | 6,144 | 752 | 800 | 7,696 | |
| Textbook | Chunk | 64 | 8 | 9 | 81 |
| Q&A Pair | 1,182 | 162 | 161 | 1,505 |
| Min | Mean | Max | |
| Question | 9 | 35 | 82 |
| Answer | 6 | 56 | 271 |
| Chunk (Ver.0) | 66 | 843 | 9,528 |
| Chunk (Ver.1) | 106 | 884 | 9,528 |
The original slide (Figure 3) often contains tables, graphs, and images, which are simplified into plain text descriptions during data processing. To evaluate performance on questions referencing these elements, we sample 143 multi-modal chunks and generated 1,861 targeted questions. Evaluation results are provided in Section 4.1.3. Additionally, we extract and prepend headers summarizing each report, including the Test Name, Region, State, and Purpose, which significantly improve the retrieval phase (See Section 4.1.2). Domain experts are responsible for the entire process, including prompt engineering and reviewing intermediate and final outputs. Detailed prompt templates for QA generation, Markdown conversion, and evaluation are available in our released repository (see Section 5).
3.2. RAG Components
We strategically select backbone models and training strategies for the Retriever & Re-ranker (Section 3.2.1) and Generation Model (Section 3.2.2), with a focus on effective integration under the constraints of on-premise deployment, domain adaptation, and Korean language support.
3.2.1. Retrieval & Re-ranking
We employ a Dual-Encoder as our retriever, fine-tuning it on our training data. For a question , the model retrieves the top chunks from chunks , based on the [CLS] token embedding similarity between and each chunk . We use BGE-M3 (Chen et al., 2024) as the backbone—a multilingual encoder capable of processing Korean—and adopt publicly available fine-tuning code 333https://github.com/FlagOpen/FlagEmbedding. The prcocess can be described as follows:
Our re-ranker is a point-wise Cross-encoder initialized with the weights of the fine-tuned retriever. It is trained on a classification task (Nogueira et al., 2019) that takes a single pair as input and outputs a scalar relevance score. It re-ranks , obtained from the retrieval stage, to extract , which will be passed to the generation model. The process is formalized as follows:
In domain-specific settings where the relevance of specific entities (e.g., car models, test types, or component names) must be emphasized, re-ranking should account for not only general semantic similarity between queries and documents but also fine-grained, term-level relevance. To optimize , we apply Token Selection (TS) + Term Control Layer (TCL) training method from RRADistill (Choi et al., 2024). It effectively integrates the importance of the general semantic and specific query terms during the training process.
TS+TCL are used during training only, while inference uses the standard Cross-Encoder only with . The label is binary: for relevant pairs and for irrelevant pairs. Negative sampling retrieves the top 10 chunks from the train data for each query , excluding the gold chunk, and randomly selects three from the rest.
3.2.2. Answer Generation
@@
We fine-tune open-source LLMs on our training dataset, which consists of pairs derived from our data generation pipeline (Section 3.1). During fine-tuning, the question and the answer are concatenated into a single sequence , and the model is trained in an auto-regressive manner to predict the next token. Specifically, we choose Qwen-2.5 444We fine-tune Qwen 2.5 (14B, 72B) on our training dataset, performing full fine-tuning on the 14B model and applying Low-Rank Adaptation (LoRA) (Hu et al., 2021) for the 72B model. (Yang et al., 2024) as our backbone LLMs, one of the few multilingual models that officially supports Korean. Notably, there are currently no Korean-centric open-source LLMs with sufficient context length to address our requirements.
3.3. Reference Matching Algorithm
@@
We propose a reference matching algorithm that enables traceability to the original source documents. The algorithm segments the generated answer and links each segment to relevant reference documents using a Dynamic Programming (DP) (Bellman, 1954) approach, leveraging our re-ranker (Section 3.2.1). Algorithm 1 outlines the detailed procedure. In Step 1, the algorithm computes the scores for all possible sentence subsequences against the top-k chunks , using re-ranker. Here, refers to candidate subsequences of sentences within considered for matching. In Step 2, it selects the optimal segment-chunk combinations, updating scores and recording the best choices for each ending sentence (e.g., if the answer consists of five sentences like , the choices might be ). Finally, in Step 3, backtracking is performed to retrieve and output the best matches from the choices (e.g., ), i.e., the first to fourth sentences of are grouped as one segment aligned to , and the fifth sentence is aligned to .
3.4. Deployment
We internally deployed the on-premise RAG system—comprising a 0.5B retriever, a 0.5B re-ranker, and a 72B LLM— for pilot testing. Employees evaluated the system and provided positive feedback. A full-scale deployment is currently being planned.
4. Experiment
4.1. Retrieval & Re-ranking
4.1.1. Evaluation
@@
We evaluate the retriever and re-ranker using Mean Average Precision (MAP@k) and Success@k (Manning et al., 2008). MAP@k measures the ranking quality by averaging precision over results up to rank , while Success@k indicates the proportion of queries with at least one relevant result in the top . The high Success@k score indicates that relevant chunks are retrieved, while the improved MAP@k highlights better ranking of those retrieved chunks. The evaluation is conducted using test set questions, but the chunk pool for retrieval included all splits (training, validation, and test) to ensure sufficient data and avoid biases from the limited test set size.
4.1.2. Result
@@
Table 4 shows that prepending a header to each chunk—adding only a small number of tokens (see Table 3)—significantly enhances retrieval performance. Fine-tuning the BGE-M3 model on Ver.1 (BGE-FT) further improves results notably, demonstrating the importance of task-specific model adaptation to optimize performance. Table 5 shows that the re-ranking phase notably enhances the precision of retrieved information. The results indicate that ranking performance is optimal when the number of retrieved chunks is limited to 10. Despite the constraints of retriever failure, this improves retrieval results by about 4% in terms of MAP@1. In addition, our ablation study confirmed that applying TS+TCL during training slightly improves ranking performance.
| Model | Train | Test | MAP@1 | MAP@5 | MAP@10 | ||
N/A | Ver.0 | 0.1978 | 0.2672 | 0.2766 | ||
| Ver.1 | 0.2991 | 0.3978 | 0.4091 | ||||
@@
| MAP@1 | MAP@5 | MAP@10 | Success@5 | |
| 10 | 0.6444 (± 0.01) | 0.7416 (± 0.01) | 0.7464 (± 0.01) | 0.8839 (± 0.00) |
| 20 | 0.6362 (± 0.01) | 0.7353 (± 0.01) | 0.7413 (± 0.01) | 0.8822 (± 0.01) |
| 30 | 0.6328 (± 0.02) | 0.7316 (± 0.02) | 0.7378 (± 0.02) | 0.8796 (± 0.01) |
4.1.3. Multi-modal Specific Questions
Table 6 summarizes the performance on questions requiring information from tables, graphs, and images. Compared to general questions (Section 4.1.2), the results are notably lower, with the retriever facing significant challenges. Even at = 30, Success@k only reaches 83.61%, underscoring a major bottleneck in the retrieval phase. Despite these challenges, the re-ranker proved effective, delivering a substantial performance improvement of approximately 15% in MAP@1 at = 20. These results underscore the need for further advancements to better handle multi-modal questions.